Your company is deciding whether to invest in a new machine. The new machine will increase cash flow by $322,000 per year. You believe the technology used in the machine has a 10-year life; in other words, no matter when you purchase the machine, it will be obsolete 10 years from today. The machine is currently priced at $1,720,000. The cost of the machine will decline by $105,000 per year until it reaches $1,195,000, where it will remain.

If your required return is 13 percent, calculate the NPV if you purchase the machine today.

If your required return is 13 percent, calculate the NPV if you wait to purchase the machine until year 6.
Answer #1

If your required return is 13 percent, calculate the NPV if you purchase the machine today.

Cost of the machine today = C0 = 1,720,000

The new machine will increase cash flow by $322,000 per year.

Hence, C = 322,000

Remaining life before obsolescence = 10 years = N

Discount rate, R = required return = 13%

PV of all future benefits = PV of an annuity "C" over 10 years at 13% discount rate

R × (1 _ (1 + R)-N) = 322,000 0.13x (1-(1+0.13)10)

=  1,747,250.40

NPV = -C0 + PV of all future benefits = -1,720,000 + 1,747,250.40 =  $ 27,250.40

If your required return is 13 percent, calculate the NPV if you wait to purchase the machine until year 6.

The cost of the machine will decline by $105,000 per year until it reaches $1,195,000, where it will remain.

hence, the cost of machine at the end of year 6 = max (1720000 - 6 x 105000, 1195000) = $ 1,195,000 = C0

No matter when you purchase the machine, it will be obsolete 10 years from today.

Remaining life before obsolescence = 10 - 6 = 4 years = N

PV of all future benefits = PV of an annuity "C" over 10 years at 13% discount rate

R × (1 _ (1 + R)-N) = 322,000 0.13 x(1-(1+0.13)4

= 957,779.77

NPV = -C0 + PV of all future benefits = -1,195,000 +  957,779.77 = - $ 237,220.23